OpenStudy (anonymous):

What is that the period, amplitude, and phase shift of f(x) = 3cos2(x- pi/4)+1 ? b) List the five critical points for the first cycle. c) Use this information to sketch the graph of f(x) for two complete cycles. Determine ¼ wave and starting point. Clearly label the scale on both axes.

OpenStudy (loser66):

any idea?

OpenStudy (anonymous):

@Loser66 no :( please teach me

OpenStudy (loser66):

:( how to teach ????

OpenStudy (loser66):

willing to study?? need patience (I need it , too) OK?

OpenStudy (loser66):

@ranga I don't know how to teach, please, help

OpenStudy (anonymous):

amplitude will be 3. period :2pi /2 = pi . phase shift is 2*pi/4 = pi/2.critical points are when first or second derivative = 0.

OpenStudy (anonymous):

@sre can you help me figure out the critical points, I haven't learned how to calculate them yet

OpenStudy (anonymous):

general form is \[f(x)=Acos(\omega t+\phi _{0})\]

OpenStudy (anonymous):

A:amplitude omega: angular frequency t: time phi_0:phase

OpenStudy (anonymous):

\[\omega=\frac{ 2\pi }{ T }\] T is the period

OpenStudy (anonymous):

I believe the phase of the function is everything within the trig function and the period can be calculated from omega

OpenStudy (anonymous):

\[x-\frac{ \pi }{ 4 }=0=\phi \]

OpenStudy (dumbcow):

is that a cos^2 function?

OpenStudy (anonymous):

@sre can you guys help me graph c) please somebody? or is @dumbcow example the sketch for f(x) for two complete cycles?